# Université de Genève, CH-1211 Genève, Switzerland

### v1

@@ 4.2 The three-twist knot
We have verified numerically that the polynomials $D_{\pm s}(q)$ have integer coefficients up to $s=20$ for all three branches. We then compute the integers $c_{\pm s}$, $s\in{\mathbb{Z}}_{>0}$, using Eqs. (3.65), (3.66), which give
$\{c^{(\rm{geom})}_{s}\}_{s\in{\mathbb{Z}}_{>0}}=$$\{1,1,-8,-15,1,64,148,49,-575,-1599,-1088,5088,\dots\}\,,$ (4.166a)
$\{c^{(\rm{conj})}_{s}\}_{s\in{\mathbb{Z}}_{>0}}=$ $(-1)\{1,5,28,101,401,1544,6077,24101,96418,388205,1571307,6387608,\dots\}\,,$ (4.166b)
$\{c^{(\rm{self})}_{s}\}_{s\in{\mathbb{Z}}_{>0}}=$$\{2,10,38,122,402,1486,5931,24058,96995,389810,1572397,6382526\dots\}\,,$ (4.166c)
@@
while $c^{(\rm{conj})}_{-s}=[s]_{2}+5[s+1]_{2}$, where $[s]_{2}$ denotes the remainder of the division of $s$ by $2$, and $c^{(\rm{self})}_{-s}=1$ for all $s\in{\mathbb{Z}}_{>0}$. We observe that the above sequences of integers satisfy the relation
$c^{(\rm{geom})}_{s}+c^{(\rm{conj})}_{s}+c^{(\rm{self})}_{s}=c^{(\rm{conj})}_{-s}+c^{(\rm{self})}_{-s}\,,\quad s\in{\mathbb{Z}}_{>0}\,.$ (4.167)
@@
which corresponds to the analogous formula in Eq. (4.146) for the figure-eight knot. Note that, in both our examples, the sum of the integer sequences $\{c_{s}\}$ over all branches equals the sum of the integer sequences $\{c_{-s}\}$, which is a periodic function of $s\in{\mathbb{Z}}_{>0}$. Again, the alternative computational approach of Appendix B agrees with the above results.
@@
Let us now go back to the quantum $A$-polynomial for the $\bm{5}_{2}$-knot written in Eq. (2.37). The corresponding $q$-difference equation in the form of Eq. (2.20) is of third order. Precisely, we have
$\hat{A}_{\bm{5}_{2}}(\hat{P},\hat{X},q)\chi^{(\alpha),\rm{WKB}}(u;\hbar)=0\,,$ (4.168)
@@
where $\chi^{(\alpha),\rm{WKB}}(u;\hbar)$ is the resummed perturbative wave function in Eq. (3.87) associated with the branch $\alpha$. In the rational case of $q={\rm e}^{2\pi{\rm i}{P\over Q}}$ with $P,Q\in{\mathbb{Z}}_{>0}$ coprime, Eq. (4.168) can be solved in closed form by directly applying the results of Section 3.4.2. As for the $\bm{4}_{1}$-knot, let us consider the case $\alpha=\rm{geom}$ and drop the label to simplify our notation. After appropriately accounting for the exponential prefactor containing the functions $s_{0,1}(x)$ in Eq. (4.158a), the $q$-difference equation above assumes the form in Eq. (3.112) with the identifications
$A(X)=q^{-2}X^{-2}\frac{C_{2}(X,q)}{C_{3}(X,q)}\,,\quad B(X)=q^{-2}X^{-4}\frac{C_{1}(X,q)}{C_{3}(X,q)}\,,\quad C(X)=X^{-6}\frac{C_{0}(X,q)}{C_{3}(X,q)}\,,$ (4.169)
@@
where $C_{0},C_{1},C_{2},C_{3}$ are the coefficient functions defined in Eq. (2.2). Substituting the expressions in Eq. (4.169) into Eq. (3.119) and computing the matrix ${\cal M}_{Q}(X)$, we can then derive explicit formulae for $\Pi\upsilon$ and $\upsilon_{0}$ from Eqs. (3.130), (3.132). We stress that the ambiguity in the definition of $\Theta_{2}$ is resolved by requiring Eq. (3.109) to be satisfied with $\alpha=\text{geom}$, which implies the choice $m=2$. Plugging these into Eq. (3.133), we find the functions $\Upsilon(X)$ and $\varphi_{Q}(X)$ for the geometric branch of the three-twist knot. We can easily use the resulting explicit expression for $\varphi_{Q}(X)$ to numerically evaluate the integer constants $c_{s}$, $s\in{\mathbb{Z}}_{>0}$, for the geometric branch by applying Eq. (3.86) for any choice of $P,Q\in{\mathbb{Z}}_{>0}$ coprime, thus producing a third independent computation which is in perfect agreement with the previous two. The same discussion can be applied to the conjugate and self-conjugate branches of the $\bm{5}_{2}$-knot, in which case the appropriate functions $s_{0,1}(x)$ are written in Eqs. (4.158b), (4.158c) and the ambiguity in the choice of $m$ in Eq. (3.131c) is fixed by imposing the condition in Eq. (3.109) with $\alpha=\text{conj},\text{self}$, yielding $m=1$ and $m=0$, respectively. Again, the corresponding integer constants $c_{\pm s}$, $s\in{\mathbb{Z}}_{>0}$, for the conjugate and self-conjugate branches match our previous results.
@@ 5 The wave function from the state integral
In this section, we consider the Andersen–Kashaev invariant ak ; ak2 ; dimofte-rev $\chi_{{\cal K}}(u;\hbar)$ in its original form as a finite-dimensional state integral, whose integrand is a product of Faddeev’s quantum dilogarithm functions faddeev . We evaluate it directly at rational values of $\hbar$ by applying the techniques of garou-kas , thus providing a third method for the computation of the exact wave function. For simplicity, we focus on the case of the figure-eight knot and show that this third approach agrees with the two previous methods presented in Section 3. A similar calculation in the case ${\mathsf{b}}^{2}=1$ was done in unpublished work by Szabolcs Zakany.
@@ 5.1 Evaluation at rational points
Following the conventions of ak , the state integral for the $\bm{4}_{1}$-knot takes the form
[…]
@@
$f(y)=\frac{\Phi_{{\mathsf{b}}}\left(\frac{x-y}{2\pi{\mathsf{b}}}\right)}{\Phi_{{\mathsf{b}}}\left(\frac{y}{2\pi{\mathsf{b}}}\right)}{\rm e}^{\frac{{\rm i}xy}{\pi{\mathsf{b}}^{2}}}\,,$ (5.171)
@@
$y$ $=x-\pi{\rm i}\left(2M+1\right)-\pi{\rm i}{\mathsf{b}}^{2}\left(2N+1\right)\,,$ (5.172a)
$y$ $=-\pi{\rm i}\left(2M+1\right)-\pi{\rm i}{\mathsf{b}}^{2}\left(2N+1\right)\,,$ (5.172b)
where $M,N\in{\mathbb{N}}$, as dictated by the singular structure of Faddeev’s quantum dilogarithm in Appedix C. Therefore, $f(y)$ is a holomorphic function in the upper half-plane. Let us now fix the rational value
${\mathsf{b}}^{2}=\frac{P}{Q}\,,\quad P,Q\in{\mathbb{Z}}_{>0}\;\;\text{coprime}\,,$ (5.173)
@@
$f(y+2\pi{\rm i}P)=f(y+2\pi{\rm i}{\mathsf{b}}^{2}Q)$ $=\Phi_{{\mathsf{b}}}\left(\frac{x-y}{2\pi{\mathsf{b}}}-{\rm i}{\mathsf{b}}Q\right)\Phi_{{\mathsf{b}}}\left(\frac{y}{2\pi{\mathsf{b}}}+{\rm i}{\mathsf{b}}Q\right)^{-1}{\rm e}^{\frac{{\rm i}xy}{\pi{\mathsf{b}}^{2}}-2xQ}$ (5.174)
$=f(y)\left(1+{\rm e}^{Q(x-y)-\pi{\rm i}PQ}\right)\left(1+{\rm e}^{Qy+\pi{\rm i}PQ}\right){\rm e}^{-2xQ}\,,$
@@
$f(y-2\pi{\rm i}P)=f(y-2\pi{\rm i}{\mathsf{b}}^{2}Q)$ $=\Phi_{{\mathsf{b}}}\left(\frac{x-y}{2\pi{\mathsf{b}}}+{\rm i}{\mathsf{b}}Q\right)\Phi_{{\mathsf{b}}}\left(\frac{y}{2\pi{\mathsf{b}}}-{\rm i}{\mathsf{b}}Q\right)^{-1}{\rm e}^{\frac{{\rm i}xy}{\pi{\mathsf{b}}^{2}}+2xQ}$ (5.175)
$=f(y)\left(1+{\rm e}^{Q(x-y)+\pi{\rm i}PQ}\right)^{-1}\left(1+{\rm e}^{Qy-\pi{\rm i}PQ}\right)^{-1}{\rm e}^{2xQ}\,,$
where we have applied the periodicity formulae for $\Phi_{{\mathsf{b}}}(z)$ at rational values in Eqs. (C.10a), (C.10b). Since $P,Q\in{\mathbb{Z}}_{>0}$ and thus ${\rm e}^{\pi{\rm i}PQ}={\rm e}^{-\pi{\rm i}PQ}$, it follows from Eqs. (5.174), (5.175) that $f(y)$ satisfies the functional equation
@@
[…]
and therefore we can use Lemma 2.1 from garou-kas to write
[…]
Let us now compute the integral on the RHS of Eq. (5.178) applying the residue theorem. We start by using Eq. (5.174) to express the integrand in the form
$\frac{f(y)}{1-f(y+2\pi{\rm i}P)/f(y)}=\frac{f(y){\rm e}^{2xQ}}{G(x,y)}\,,$ (5.179)
@@
$G(x,y)={\rm e}^{2xQ}-\left(1+{\rm e}^{Q(x-y)-\pi{\rm i}PQ}\right)\left(1+{\rm e}^{Qy+\pi{\rm i}PQ}\right)\,.$ (5.180)
Since the numerator in Eq. (5.179) is holomorphic in the upper half of the complex plane, the relevant singularities of the integrand are the solutions to $G(x,y)=0$ at fixed $x\in{\mathbb{R}}$ which lie within the domain of integration. Specifically, these are
$y_{\pm,k}(x)=2\pi{\rm i}\frac{k}{Q}-{\rm i}\pi P-\frac{1}{Q}\log\left(-\frac{1+{\rm e}^{xQ}-{\rm e}^{2xQ}\pm\sqrt{\Delta({\rm e}^{xQ})}}{2{\rm e}^{xQ}}\right)\,,$ (5.181)
where $k\in{\mathbb{Z}}$ such that $10\leq k\leq PQ-$ 1, the function $\Delta(X)$ is defined by

### v2

@@ 4.2 The three-twist knot
We have verified numerically that the polynomials $D_{\pm s}(q)$ have integer coefficients up to $s=20$ for all three branches. We then compute the integers $c^{(\alpha)}_{\pm s}$, $s\in{\mathbb{Z}}_{>0}$, using equations (3.27) and (3.28), which give
$\big\{c^{({\rm geom})}_{s}\big\}_{s\in{\mathbb{Z}}_{>0}}=\{1,1,-8,-15,1,64,148,49,-575,-1599,-1088,5088,\dots\},$ (4.33a)
$\big\{c^{({\rm conj})}_{s}\big\}_{s\in{\mathbb{Z}}_{>0}}=(-1)\{1,5,28,101,401,1544,6077,24101,96418,388205,1571307,$
[…]
$\big\{c^{({\rm self})}_{s}\big\}_{s\in{\mathbb{Z}}_{>0}}=\{2,10,38,122,402,1486,5931,24058,96995,389810,1572397,$
[…]
@@
while $c^{({\rm conj})}_{-s}=[s]_{2}+5[s+1]_{2}$, where $s]_{2}$ denotes the remainder of the division of $s$ by $2$, and $c^{({\rm self})}_{-s}=1$ for all $s\in{\mathbb{Z}}_{>0}$. We observe that the above sequences of integers satisfy the relation
$c^{({\rm geom})}_{s}+c^{({\rm conj})}_{s}+c^{({\rm self})}_{s}=c^{({\rm conj})}_{-s}+c^{({\rm self})}_{-s},\qquad s\in{\mathbb{Z}}_{>0},$ (4.34)
corresponding to the analogous formula in equation (4.13) for the figure-eight knot. Note that, in both our examples, the sum over all non-abelian branches of the integers $c^{(\alpha)}_{s}$ at fixed $s$ equals the sum of the integers $c^{(\alpha)}_{-s}$, that is,
$\sum_{\alpha}c^{(\alpha)}_{s}=\sum_{\alpha}c^{(\alpha)}_{-s},\qquad s\in{\mathbb{Z}}_{>0},$ (4.35)
@@
which appears to be a periodic function of $s$. Again, the alternative computational approach of Appendix B agrees with the above results.
@@
Let us now go back to the quantum $A$-polynomial for the $\bm{5}_{2}$-knot written in equation (2.38). The corresponding $q$-difference equation in the form of equation (2.21) is of third order. Precisely, we have
$\hat{A}_{\bm{5}_{2}}\bigl(\hat{P},\hat{X},q\bigr)\chi^{(\alpha),{\rm WKB}}(u;\hbar)=0,$ (4.36)
@@
where $\chi^{(\alpha),{\rm WKB}}(u;\hbar)$ is the resummed perturbative wave function in equation (3.49) associated with the branch $\alpha$. In the rational case of $q={\rm e}^{2\pi{\rm i}{P\over Q}}$ with $P,Q\in{\mathbb{Z}}_{>0}$ coprime, equation (4.36) can be solved in closed form by directly applying the results of Section 3.4.2. As for the $\bm{4}_{1}$-knot, let us consider the case $\alpha={\rm geom}$ and drop the label to simplify our notation. After appropriately accounting for the exponential prefactor containing the functions $s_{0,1}(x)$ in equation (4.25a), the $q$-difference equation above assumes the form in equation (3.74) with the identifications
$\begin{split}&A(X)=q^{-2}X^{-2}{{C_{2}(X,q)\over C_{3}(X,q)}},\qquad B(X)=q^{-2}X^{-4}{{C_{1}(X,q)\over C_{3}(X,q)}},\\ &C(X)=X^{-6}{{C_{0}(X,q)\over C_{3}(X,q)}},\end{split}$ (4.37)
@@
where $C_{0}$, $C_{1}$, $C_{2}$, $C_{3}$ are the coefficient functions defined in equation (2.39). Substituting the expressions in equation (4.37) into equation (3.81) and computing the matrix ${\cal M}_{Q}(X)$, we can then derive explicit formulae for $\Pi\upsilon$ and $\upsilon_{0}$ from equations (3.92) and (3.94). We stress that the ambiguity in the definition of $\Theta_{2}$ is resolved by requiring equation (3.71) to be satisfied with $\alpha={\rm geom}$}, which implies the choice $m=2$. Plugging these into equation (3.95), we find the functions $\Upsilon(X)$ and $\varphi_{Q}(X)$ for the geometric branch of the three-twist knot. We can easily use the resulting explicit expression for $\varphi_{Q}(X)$ to numerically evaluate the integer constants $c_{s}$, $s\in{\mathbb{Z}}_{>0}$, for the geometric branch by applying equation (3.48) for any choice of $P,Q\in{\mathbb{Z}}_{>0}$ coprime, thus producing a third independent computation which is in perfect agreement with the previous two. The same discussion can be applied to the conjugate and self-conjugate branches of the $\bm{5}_{2}$-knot, in which case the appropriate functions $s_{0,1}(x)$ are written in equations (4.25b) and (4.25c) and the ambiguity in the choice of $m$ in equation (3.93c) is fixed by imposing the condition in equation (3.71) with $\alpha=\mathrm{conj},\mathrm{self}$, yielding $m=1$ and $m=0$, respectively. Again, the corresponding integer constants $c_{\pm s}$, $s\in{\mathbb{Z}}_{>0}$, for the conjugate and self-conjugate branches match our previous computations.
4.3 Proofs of admissibility
As mentioned in Section 3.1, the conjectured integrality property of the resummed perturbative wave function in equation (3.1) can be stated as admissibility in the sense of Kontsevich–Soibelman.
Recall that a formal power series in one variable $z$ with coefficients in the ring of formal Laurent series ${\mathbb{Z}}\bigl(\!\bigl(q^{1/2}\bigr)\!\bigr)$, that is, $F(z;q)=\sum_{n\geq 0}A_{n}z^{n}\in{\mathbb{Z}}\bigl(\!\bigl(q^{1/2}\bigr)\!\bigr)[\![z]\!]$, is admissible if and only if it can be expressed as a (possibly infinite) product of quantum dilogarithms of the form
$F(z;q)=\prod_{s\geq 1}\prod_{i\in{\mathbb{Z}}}\bigl(q^{i/2}z^{s};q\bigr)_{\infty}^{c(s,i)}\in 1+z{\mathbb{Z}}\bigl(\!\bigl(q^{1/2}\bigr)\!\bigr)[\![z]\!],\qquad c(s,i)\in{\mathbb{Z}},$ (4.38)
where $c(s,i)=0$ for $|i|\gg 0$ and fixed $s$. It follows that admissible series form a group under multiplication. Note that the formula above is equivalent to
$F(z;q)=\exp\left(\sum_{k,s\geq 1}{{f_{s}\bigl(q^{k/2}\bigr)\over k\bigl(q^{k}-1\bigr)}}z^{sk}\right),\qquad f_{s}(q)\in{\mathbb{Z}}\big[q^{\pm}\big],$ (4.39)
which reproduces in $z$ the same integrality structure we conjectured for the $X$, $X^{-1}$-series in equation (3.6) with one caveat. More precisely, for all choices of non-abelian branch $\alpha$ and up to the exponential prefactors, the resummed WKB wave function in equation (3.1) is expected to be a product of two admissible series in $X$ and $X^{-1}$, respectively, such that $q^{-k\sigma_{s}}D_{s}\bigl(q^{k}\bigr)=f_{s}\bigl(q^{k/2}\bigr)\in{\mathbb{Z}}\big[q^{\pm}\big]$ for all $k$, $s$.
The notion of admissibility can be appropriately extended [^16] In fact, admissibility is defined in [34] for series in $R\bigl(\!\bigl(q^{1/2}\bigr)\!\bigr)[\![\{z_{i}\}_{i\in I}]\!]$, where $R$ is a $\lambda$-ring and $q^{1/2},\{z_{i}\}_{i\in I}$ are line elements. to apply to multi-variable series in ${\mathbb{Z}}\bigl(\!\bigl(q^{1/2}\bigr)\!\bigr)[\![\{z_{i}\}_{i\in I}]\!]$, where $I={1,\dots,|I|}$. The authors of [34] proved that admissibility is preserved under certain transformations. In particular, for any given symmetric integer matrix $B=(b_{i,j})_{i,j=1,\dots,|I|}$, a series
$F(z;q)=\sum_{\{n_{i}\}\in{\mathbb{Z}}_{\geq 0}^{|I|}}A_{n}\prod_{i\in I}z_{i}^{n_{i}}\in{\mathbb{Z}}\bigl(\!\bigl(q^{1/2}\bigr)\!\bigr)[\![\{z_{i}\}_{i\in I}]\!]$ (4.40)
is admissible if and only if the twisted series
$\tilde{F}(z;q)=\sum_{\{n_{i}\}\in{\mathbb{Z}}_{\geq 0}^{|I|}}\bigl(-q^{1/2}\bigr)^{\sum_{i,j=1}^{|I|}b_{i,j}n_{i}n_{j}}A_{n}\prod_{i\in I}z_{i}^{n_{i}}$ (4.41)
is admissible [34, Theorem 9]. In what follows, we will refer to this statement as the Kontsevich–Soibelman (KS) theorem.
We will now show that the $q$-special functions appearing as $\exp\bigl(\phi^{(\alpha),{\rm WKB}}(X;q)\bigr)$ in the holomorphic/antiholomorphic block decompositions in equations (4.1) and (4.21) for the non-perturbative wave functions of the figure-eight and three-twist knots are admissible. [^17] We thank the anonymous referee for providing us with a sketch of the proof.
Proof of admissibility for the $\bm{4}_{1}$-knot..
Let us consider each classical branch separately.
(1) For $\alpha={\rm geom}$, equations (4.2) and (4.3a) imply the $q$-series representation
${\rm e}^{\phi^{({\rm geom}),{\rm WKB}}(X;q)}$ $=(qX;q)_{\infty}\sum_{n=0}^{\infty}{{(-1)^{n}q^{n(n+1)/2}\over(q;q)_{n}(qX;q)_{n}}}X^{2n}$ (4.42)
$=\sum_{m,n=0}^{\infty}q^{nm}{{(-1)^{n}q^{n(n+1)/2}\over(q;q)_{n}}}X^{2n}{{(-1)^{m}q^{m(m+1)/2}\over(q;q)_{m}}}X^{m},$
where we have applied the identities in equations (C.3) and (C.5a) in the second equality. Observe that the twist factor $q^{nm}$ corresponds to the symmetric integer matrix
$B_{\bm{4}_{1}}=\begin{pmatrix}0&1\\ 1&0\end{pmatrix}$, (4.43)
while the untwisted series
$\sum_{m,n=0}^{\infty}{{(-1)^{n}q^{n(n+1)/2}\over(q;q)_{n}}}X^{2n}{{(-1)^{m}q^{m(m+1)/2}\over(q;q)_{m}}}X^{m}=\bigl(qX^{2};q\bigr)_{\infty}(qX;q)_{\infty}$ (4.44)
is the product of two quantum dilogarithms and therefore admissible. We conclude by means of the KS theorem.
(2) For $\alpha={\rm conj}$, equations (4.2) and (4.3b) imply the $q$-series representation
${\rm e}^{\phi^{({\rm conj}),{\rm WKB}}(X;q)}$ $=\bigl(qX^{-1};q\bigr)_{\infty}\sum_{n=0}^{\infty}{{(-1)^{n}q^{n(n+1)/2}\over(q;q)_{n}\bigl(qX^{-1};q\bigr)_{n}}}X^{n}$
$=\theta(X;q)\sum_{m,n=0}^{\infty}q^{-nm}{{X^{2n}\over(q;q)_{n}}}{{X^{m}\over(q;q)_{m}}},$ (4.45)
where we have applied the identities in equations (C.2b), (C.3), and (C.5b) in the second equality, and we have denoted by
$\theta(X;q)=\bigl(qX^{-1};q\bigr)_{\infty}(X;q)_{\infty}$ (4.46)
the $q$-theta function. Observe that the twist factor $q^{-nm}$ corresponds to the symmetric integer matrix
$B_{\bm{4}_{1}}=\begin{pmatrix}0&-1\\ -1&0\end{pmatrix}$, (4.47)
while the untwisted series
$\sum_{m,n=0}^{\infty}{{X^{2n}\over(q;q)_{n}}}{{X^{m}\over(q;q)_{m}}}={{1\over\bigl(X^{2};q\bigr)_{\infty}(X;q)_{\infty}}}$ (4.48)
is the product of two quantum dilogarithms and therefore admissible. Again, we conclude by means of the KS theorem. ∎
Proof of admissibility for the $\bm{5}_{2}$-knot..
Let us consider each classical branch separately.
(1) For $\alpha={\rm geom}$, equations (4.22) and (4.23a) imply that
${\rm e}^{\phi^{({\rm geom}),{\rm WKB}}(X;q)}$
$\qquad=(qX;q)_{\infty}\bigl(qX^{2};q\bigr)_{\infty}\sum_{n=0}^{\infty}{{q^{n(n+1)}\over(q;q)_{n}(qX;q)_{n}\bigl(qX^{2};q\bigr)_{n}}}X^{2n}$
$\qquad=\sum_{m,n,k=0}^{\infty}q^{nm+nk}{{q^{n(n+1)}\over(q;q)_{n}}}X^{2n}{{(-1)^{m}q^{m(m+1)/2}\over(q;q)_{m}}}X^{m}{{(-1)^{k}q^{k(k+1)/2}\over(q;q)_{k}}}X^{2k},$ (4.49)
where we have used the identities in equations (C.3) and (C.5a) in the second step. Observe that the twist factor $q^{nm+nk}$ corresponds to the symmetric integer matrix
$B_{\bm{5}_{2}}=\begin{pmatrix}0&1&1\\ 1&0&0\\ 1&0&0\\ \end{pmatrix}$, (4.50)
while the untwisted series is the product of three admissible series [^18] The third factor in equation (4.51) can be thought of as the twist of $1/(qX;q)_{\infty}$ by a factor $q^{n^{2}}$, which gives a one-dimensional matrix $B=(2)$.
$\bigl(qX^{2};q\bigr)_{\infty}(qX;q)_{\infty}\sum_{n=0}^{\infty}{{q^{n(n+1)}\over(q;q)_{n}}}X^{2n}.$ (4.51)
The KS theorem applies.
(2) For $\alpha={\rm conj}$, equations (4.22) and (4.23b) imply that
${\rm e}^{\phi^{({\rm conj}),{\rm WKB}}(X;q)}$ $=\bigl(qX^{-1};q\bigr)_{\infty}\bigl(qX^{-2};q\bigr)_{\infty}\sum_{n=0}^{\infty}{{q^{n(n+1)}\over(q;q)_{n}\bigl(qX^{-1};q\bigr)_{n}\bigl(qX^{-2};q\bigr)_{n}}}X^{-2n}$ (4.52)
$=\theta(X;q)\theta\bigl(X^{2};q\bigr)\sum_{m,n,k=0}^{\infty}q^{-nm-nk}{{X^{2n}\over(q;q)_{n}}}{{X^{m}\over(q;q)_{m}}}{{X^{2k}\over(q;q)_{k}}},$
where we have used the identities in equations (C.2b), (C.3), and (C.5b) in the second step. The $q$-theta function $\theta(X;q)$ is introduced in equation (4.46). Observe that the twist factor $q^{-nm-nk}$ corresponds to the symmetric integer matrix
$B_{\bm{5}_{2}}=\begin{pmatrix}0&-1&-1\\ -1&0&0\\ -1&0&0\\ \end{pmatrix}$, (4.53)
while the untwisted series is the product of three quantum dilogarithms
$1\over\bigl(X^{2};q\bigr)^{2}_{\infty}(X;q)_{\infty}$}}. (4.54)
The KS theorem applies.
(3) For $\alpha={\rm self}$, equations (4.22) and (4.23c) imply that
${\rm e}^{\phi^{({\rm self}),{\rm WKB}}(X;q)}$ $=(qX;q)_{\infty}\bigl(qX^{-1};q\bigr)_{\infty}\sum_{n=0}^{\infty}{{q^{n(n+1)}\over(q;q)_{n}(qX;q)_{n}\bigl(qX^{-1};q\bigr)_{n}}}$
$=\theta(X;q)\sum_{m,n,k=0}^{\infty}q^{nm-nk}{{(-1)^{n}q^{n(n+1)/2}\over(q;q)_{n}}}{{(-1)^{m}q^{m(m+1)/2}\over(q;q)_{m}}}X^{m}{{X^{k}\over(q;q)_{k}}},$ (4.55)
where we have used the identities in equations (C.2b), (C.3), (C.5a), and (C.5b) in the second step. The $q$-theta function $\theta(X;q)$ is introduced in equation (4.46). Observe that the twist factor $q^{nm-nk}$ corresponds to the symmetric integer matrix
$B_{\bm{5}_{2}}^{\prime}=\begin{pmatrix}0&1&-1\\ 1&0&0\\ -1&0&0\\ \end{pmatrix}$, (4.56)
while the untwisted series is the product of three quantum dilogarithms
$(q;q)_{\infty}(qX;q)_{\infty}{{1\over(X;q)_{\infty}}}.$ (4.57)
Once more, the KS theorem applies. ∎
We conclude by noting that the admissible series in $X^{-1}$ are remarkably simpler than the corresponding $X$-series in the factorisations in equations (4.45), (4.52), and (4.55). Particularly, they appear within products of $q$-theta functions in both examples of the $\bm{4}_{1}$- and $\bm{5}_{2}$-knots.
@@ 5 The wave function from the state integral
In this section, we consider the Andersen–Kashaev invariant $\chi_{{\cal K}}(u;\hbar)$ in its original form as a finite-dimensional state integral [2, 3, 9], whose integrand is a product of Faddeev’s quantum dilogarithm functions [13]. We evaluate it directly at rational values of $\hbar$ by applying the techniques of [18], thus providing a third method for the computation of the exact wave function. For simplicity, we focus on the case of the figure-eight knot and show that this third approach agrees with the two previous methods presented in Section 3. A similar calculation in the case ${\mathsf{b}}^{2}=1$} was done in unpublished work by Szabolcs Zakany.
@@ 5.1 Evaluation at rational points
Following the conventions of [2], the state integral for the $\bm{4}_{1}$-knot takes the form [^19] Note that the expressions in (4.1) and (5.1) differ by the global exponential factor ${\rm e}^{-{{\pi{\rm i}\over 12}}({\mathsf{b}}^{2}+{\mathsf{b}}^{-2})-{{x\over 2}}-{{x\over 2{\mathsf{b}}^{2}}}}$.
$\chi_{\bm{4}_{1}}(u;\hbar)={{1\over 2\pi{\mathsf{b}}}}{\rm e}^{-{{x\over 2}}-{{x\over 2{\mathsf{b}}^{2}}}-{{{\rm i}x^{2}\over 2\pi{\mathsf{b}}^{2}}}}\int_{{\mathbb{R}}+{\rm i}0}{{\Phi_{{\mathsf{b}}}\left({{x-y\over 2\pi{\mathsf{b}}}}\right)\over\Phi_{{\mathsf{b}}}\left({{y\over 2\pi{\mathsf{b}}}}\right)}}{\rm e}^{{{{\rm i}xy\over\pi{\mathsf{b}}^{2}}}}{\rm d}y,$ (5.1)
@@
$f(y)={{\Phi_{{\mathsf{b}}}\left({{x-y\over 2\pi{\mathsf{b}}}}\right)\over\Phi_{{\mathsf{b}}}\left({{y\over 2\pi{\mathsf{b}}}}\right)}}{\rm e}^{{{{\rm i}xy\over\pi{\mathsf{b}}^{2}}}},$ (5.2)
@@
$y=x-\pi{\rm i}(2M+1)-\pi{\rm i}{\mathsf{b}}^{2}(2N+1),$ (5.3a)
$y=-\pi{\rm i}(2M+1)-\pi{\rm i}{\mathsf{b}}^{2}(2N+1),$ (5.3b)
where $M,N\in{\mathbb{N}}$, as dictated by the singularities and the zeroes of Faddeev’s quantum dilogarithm in Appendix C. Therefore, $f(y)$ is a holomorphic function in the upper half-plane. Let us now fix the rational value
${\mathsf{b}}^{2}={{P\over Q}},\qquad P,Q\in{\mathbb{Z}}_{>0}\ {\rm coprime},$ (5.4)
@@
$f(y+2\pi{\rm i}P)=f\bigl(y+2\pi{\rm i}{\mathsf{b}}^{2}Q\bigr)$ $=\Phi_{{\mathsf{b}}}\left({{x-y\over 2\pi{\mathsf{b}}}}-{\rm i}{\mathsf{b}}Q\right)\Phi_{{\mathsf{b}}}\left({{y\over 2\pi{\mathsf{b}}}}+{\rm i}{\mathsf{b}}Q\right)^{-1}{\rm e}^{{{{\rm i}xy\over\pi{\mathsf{b}}^{2}}}-2xQ}$
$=f(y)\bigl(1+{\rm e}^{Q(x-y)-\pi{\rm i}PQ}\bigr)\bigl(1+{\rm e}^{Qy+\pi{\rm i}PQ}\bigr){\rm e}^{-2xQ},$ (5.5)
@@
$f(y-2\pi{\rm i}P)$ $=f\bigl(y-2\pi{\rm i}{\mathsf{b}}^{2}Q\bigr)=\Phi_{{\mathsf{b}}}\left({{x-y\over 2\pi{\mathsf{b}}}}+{\rm i}{\mathsf{b}}Q\right)\Phi_{{\mathsf{b}}}\left({{y\over 2\pi{\mathsf{b}}}}-{\rm i}{\mathsf{b}}Q\right)^{-1}{\rm e}^{{{{\rm i}xy\over\pi{\mathsf{b}}^{2}}}+2xQ}$
$=f(y)\bigl(1+{\rm e}^{Q(x-y)+\pi{\rm i}PQ}\bigr)^{-1}\bigl(1+{\rm e}^{Qy-\pi{\rm i}PQ}\bigr)^{-1}{\rm e}^{2xQ},$ (5.6)
where we have applied the periodicity formulae for $\Phi_{{\mathsf{b}}}(z)$ at rational values in equations (C.17a) and (C.17b). Since $P,Q\in{\mathbb{Z}}_{>0}$ and thus ${\rm e}^{\pi{\rm i}PQ}={\rm e}^{-\pi{\rm i}PQ}$, it follows from equations (5.5) and (5.6) that $f(y)$ satisfies the functional equation
@@
$f(y)(f(y)-f(y+2\pi{\rm i}P))\neq 0,\qquad\forall y\in{\mathbb{R}}+{\rm i}0,$ (5.8)
and therefore we can use [18, Lemma 2.1] to write
$\int_{{\mathbb{R}}+{\rm i}0}f(y){\rm d}y=\left(\int_{{\mathbb{R}}+{\rm i}0}-\int_{{\mathbb{R}}+{\rm i}0+2\pi{\rm i}P}\right){{f(y)\over 1-f(y+2\pi{\rm i}P)/f(y)}}{\rm d}y.$ (5.9)
Let us now compute the integral on the right-hand side of equation (5.9) applying the residue theorem. We start by using equation (5.5) to express the integrand in the form
${{f(y)\over 1-f(y+2\pi{\rm i}P)/f(y)}}={{f(y){\rm e}^{2xQ}\over G(x,y)}},$ (5.10)
@@
$G(x,y)={\rm e}^{2xQ}-\bigl(1+{\rm e}^{Q(x-y)-\pi{\rm i}PQ}\bigr)\bigl(1+{\rm e}^{Qy+\pi{\rm i}PQ}\bigr).$ (5.11)
Since the numerator in equation (5.10) is holomorphic in the upper half of the complex plane, the relevant singularities of the integrand are the solutions to $G(x,y)=0$ at fixed $x\in{\mathbb{R}}$ that lie within the domain of integration. Specifically, these are
$y_{\pm,k}(x)=2\pi{\rm i}{{k\over Q}}-\pi{\rm i}P-{{1\over Q}}\log\left(-{{1+{\rm e}^{xQ}-{\rm e}^{2xQ}\pm\sqrt{\Delta\bigl({\rm e}^{xQ}\bigr)}\over 2{\rm e}^{xQ}}}\right),$ (5.12)
where $k\in{\mathbb{Z}}$ such that $0\leq k\leq PQ-1$, the function $\Delta(X)$ is defined by